OpenStudy (anonymous):

what is the answer to 4^(3x+3)=12

OpenStudy (anonymous):

4^3x.4^3=12...then we can simplify!!

OpenStudy (dumbcow):

take log of both sides \[\log 4^{3x+3} = \log 12\] \[3x+3 = \frac{\log 12}{\log 4}\] \[x = \frac{\log 12}{3\log 4} -1\]